OpenStudy (anonymous):

csc2(π2−x)1+tan2(x) simplify

OpenStudy (mathmath333):

\(\Large\rm \begin{align} \color{black}{\text{is it this}\\\csc2(2\pi−x)+tan2x}\end{align}\)

OpenStudy (noelgreco):

I'm guessing \[\csc ^{2}(\frac{ \pi }{ 2 }-x)+\tan ^{2}x\] but why should we have to guess?